what percent of the questions on the entire test did the students answer correctly.
algebra 17/20
trigonometry 11/15
geometry 20/25
I added correct with total and got 48/60. I dont know what to do next to get the percentage of correct answered. Please help. Thank you.

Answer by richard1234(7193)

Divide 48 by 60, then convert to a percent.

48/60 = 4/5 = .8 = 80%